All of these sets of quantum numbers are permissible except
n                       l                        ml                      ms

2                       2                       0                       -1/2

3                       2                       -1                      +1/2

3 1                       1                       -1/2

1                       0                       0                       +1/2
Answer #1

- principal quantum number co has the range from 11,2,...0 » Azimichal quantim numbex (I) has the range from co to (n-1)) meOn=2 l =2, mezoi ms -- 1 2 Cincorrect) n=2 l= (0) 240 oto (2-1) = 0 to 1 me = 0,1 (d) to (tl) = (0) to (to) (+1) or (t/ Ms. m0 t. I ginal l. =: 0 ou ms me=OL 2 n=1 o to .(n-1) to (lt) => a me all tolte o to o ms 2 O F (paus)) T4 -1/ Hence except ist